OpenStudy (samigupta8):

for a<0 determine all solution of the equation x^2-2aIx-aI-3a^2=0 I I symbol is modulus sign

OpenStudy (samigupta8):

@divu.mkr pls..hlp me in dis

OpenStudy (anonymous):

\[x ^{2}-2a \left| x-a \right|-3a ^{2}=0\]

OpenStudy (anonymous):

consider two cases x-a>0 x-a<0..btw you know how to solve mods?

OpenStudy (anonymous):

keep in mind if x-a<0,x<a i.e.,x<0

OpenStudy (amistre64):

consider: \[x ^{2}-2a(x-a)-3a ^{2}=0\]and \[x ^{2}-2a (a-x)-3a ^{2}=0\]

OpenStudy (samigupta8):

yaa i m solving it @amistre64

OpenStudy (samigupta8):

i m getting 2 values of x as a+-root 2a

OpenStudy (amistre64):

\[x ^{2}-2a(x-a)-3a ^{2}=0\to x^2-2x-a^2=0\] \[x = 1\pm\sqrt{1+a^2}\] \[x^{2}-2a(a-x)-3a ^{2}=0\to x^2+2x-5a^2=0\] \[x = -1\pm\sqrt{1+5a^2}\]

OpenStudy (amistre64):

plug them back into the original setup to see if they all work :)

OpenStudy (samigupta8):

wich original setup are u talking abt

OpenStudy (amistre64):

at the startof your post you were considering:\[x^2-2a|x-a|-3a^2=0\]

OpenStudy (samigupta8):

yaa right so do u want me to put the values in dis one

OpenStudy (amistre64):

of course, it is the only one that we actually care about. Splitting it allows us to find critical values to test out.

OpenStudy (samigupta8):

donn't u think it wud be tedious process

OpenStudy (amistre64):

lol, math isnt all speed and quickness

OpenStudy (amistre64):

if theres another way to test them out .. then by all means give it a shot :)

OpenStudy (amistre64):

im not getting a general result, but only specific cases for a value of a; like a=0, 1, 2sqrt2-3

OpenStudy (samigupta8):

answer is (1-root 2)a,(-1+root 6)a

OpenStudy (amistre64):

hmm, accursed moduluses :)

OpenStudy (amistre64):

-2ax ... i dropped an a :/

OpenStudy (amistre64):

\[x ^{2}-2a(x-a)-3a ^{2}=0\]\[ x^2-2ax-a^2=0\]and +2ax for the other setup then work it thru ....

OpenStudy (anonymous):

@samigupta8 ur hot

OpenStudy (samigupta8):

@RadEn pls..hlp

OpenStudy (anonymous):

oh, it's what amistre said \[x^2-2a(x-a)-3a^2=0 \hspace{55px}x^2-2a(a-x)-3a^2=0\] \[ x^2-2ax + 2a^2-3a^2=0 \hspace{55px}x^2-2a^2+2ax-3a^2=0\] \[ x^2-2ax-a^2=0 \hspace{95px}x^2+2ax-5a^2=0\] \[x=\frac{2a±\sqrt{4a^2+4a^2}}{2}\hspace{85px}x=\frac{-2a±\sqrt{4a^2+20a^2}}{2}\] \[x=a±a\sqrt{2} \hspace{115px} x= -a ± a \sqrt{6} \]

OpenStudy (samigupta8):

i got that values alraedy

OpenStudy (samigupta8):

the after process is remaining

OpenStudy (anonymous):

those are the answers though/...

OpenStudy (anonymous):

I don't follow....

OpenStudy (anonymous):

plggin them back in ?

OpenStudy (samigupta8):

yaa i m doing dat...

OpenStudy (anonymous):

hmph.

OpenStudy (samigupta8):

\[x^2-2a \left| x-a \right|-3a^2=0\] putting value as \[a+\sqrt{2}a\] \[(a+\sqrt{2}a)^2-2a \left| a+\sqrt{2} a-a\right|-3a^2\]=0 \[a^2+2a^2+2\sqrt{2}a^2-2a(\sqrt{2}a)-3a^2\]=0 llhs=rhs=0

hartnn (hartnn):

since a<0 |a| = -a

hartnn (hartnn):

|a+root 2a-a| = |root 2 a| = -root 2a

hartnn (hartnn):

-2a|a+root 2a-a| = + 2root 2 a^2 not -

OpenStudy (samigupta8):

okk...that means it wud be 4root 2 a^2 and not 0

hartnn (hartnn):

yeah, now you will get 0 for (1-root 2) a try it..

OpenStudy (samigupta8):

\[(a-\sqrt{2}a)^2-2a \left| a-\sqrt{2}a-a \right|-3a^2\] \[a^2+2a^2-2\sqrt{2}a^2-2a(-\sqrt{2}a)-3a^2\]

OpenStudy (samigupta8):

it wud give me 0=0

OpenStudy (samigupta8):

and for -a-aroot6 \[(-a-a \sqrt{6})^2-2a \left| -a(2+\sqrt{6}) \right|-3a^2\]

OpenStudy (samigupta8):

how to solve dis now shall i consider - sign or not

hartnn (hartnn):

|-a*(...)| = -a (...)

hartnn (hartnn):

|-a (2+sqrt 6)| = (2+sqrt 6) |-a| = (2+sqrt 6)(-a)\ as a<0

OpenStudy (samigupta8):

yeh dekho galat aa rha hai abb

hartnn (hartnn):

it should not come =0 and its not coming, right ?

OpenStudy (samigupta8):

yaaaaaaaaaaa

OpenStudy (samigupta8):

yeh toh 8a^2+4a^2root6 aa rha hai

hartnn (hartnn):

so, -a-aroot6 is not the solution try other it'll comes as 0

OpenStudy (samigupta8):

arrey maine wohi to try krke dekha tha

OpenStudy (samigupta8):

i hve considered x=-a-root 6 here

hartnn (hartnn):

yes, and it did not =0 , right ? so -a-aroot6 is not the solution

OpenStudy (samigupta8):

bt why it's given in answer

hartnn (hartnn):

not actually.... ,(-1+root 6)a means -a + root 6 a which is given in answer as u said

OpenStudy (samigupta8):

yaa u r crrect...sry for my mistake i thought the minus sign was outside the bracket

hartnn (hartnn):

no problem! glad you finally solved it :)